Operation Summary for February 5, 2010 - February 8, 2010 |
The Integrated Luminosity for the period from 2/1/10 to 2/8/10 was 50.47 pb-1. Friday Feb. 5 The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (57.5E10), with the Antiproton Source stacking (9.4E10), with the Linac, Booster, and the Main Injector operating normally, with the NuMI and MiniBooNE experiments taking beam, with the MTest experiment off due to a problem with a Meson M01 area LCW pump, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store. During the day shift, Operations established store 7581 at 8:03 AM with an Average Initial Luminosity (AIL) of 269.2E30. Operations established beam to MTest at 8:30 AM. During the evening shift, Operations disabled the MTest beam at 8 PM. Operations terminated store 7581 at 10:57 PM and put the TeV into shot setup. Saturday Feb. 6 The midnight shift began with TeV in shot setup, with Operations monitoring the stash (12.2E10), with Pbar stacking (49.4E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established store 7582 at 12:39 AM with an AIL of 276.5E30. Operations established beam to MTest at 4:22 AM. During the day shift, Operations terminated store 7582 at 2:28 PM and put the TeV into shot setup. During the evening shift, Operations established store 7583 at 4:07 PM with an AIL of 282.84E30. Operations disabled the MTest beam at 8 PM. Sunday Feb. 7 The midnight shift began with Operations monitoring store 7583 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7583 at 4:12 AM and put the TeV into shot setup. Operations established store 7584 at 6 AM with an AIL of 278.5E30. During the day shift, The Pelletron experienced a full discharge around 9:50 AM. It had recovered and was back online by 11:30 AM. During the evening shift, Operations terminated store 7584 at 6:09 PM, however the ramp locked at 90 GeV and required some time to correct. Operations put the TeV into shot setup and established store 7586 at 9:06 PM with an AIL of 283.8E30. Monday Feb. 8 The midnight shift began with Operations monitoring store 7586 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, a problem was found with a lead on a TeV dipole magnet (T:D1TP6D), which would require a TeV A-E access to repair. Operations terminated store 7586 at 7:33 AM. The TeV access was underway by the end of shift. The Plans for Monday The plans for today are to fix the TeV magnet lead and then resume stacking, stashing, and storing. |
|
More Information Comments and Suggestions |
| last modified 02/08/2010 email Fermilab |
| Security, Privacy, Legal | |